OpenStudy (mendicant_bias):

Don't necessarily need help, just working through a problem and posting as I go both because of the convenience of the LaTeX editor and asking for help and keeping a record.

OpenStudy (mendicant_bias):

@Concentrationalizing , how did you do the Laplace L? I thought you had to turn on packages/do things you could do only in a full-fledged LaTeX editor in order to do that.

OpenStudy (mendicant_bias):

Oh, nvm, siths did it, my bad

OpenStudy (anonymous):

Sith did it, I didn't. I have no idea how he did the proper L for laplace, haha

OpenStudy (amistre64):

\[L\{f(x)\}\] i just use an L

OpenStudy (mendicant_bias):

I do, too, just wanted to figure out how he was doing that, heh. Oh well, I'll ask him the next time he's on, for now will work through it.

OpenStudy (mendicant_bias):

\[L\left\{t\ U(t-2)\right\}; \ \ \ f(t-2)=t.\]For this abstracted example, a Y(s) doesn't necessarily exist, but I'm going to pretend like it does for convenience' sake(?)

OpenStudy (anonymous):

You just have to manually shift your function is all. Can't do anything unless it's \((t-2)\)

OpenStudy (mendicant_bias):

\[Y(s)=L\left\{t \ U(t-2)\right\};\ \ \ Y(s)-2L\left\{U(t-2)\right\}=L\left\{(t \ U(t-2)\right\}-2L\left\{U(t-2)\right\}\]

OpenStudy (mendicant_bias):

\[Y(s)=L\left\{(t-2)U(t-2)\right\}+2L\left\{U(t-2)\right\}\]

OpenStudy (mendicant_bias):

\[=\frac{e^-2s}{s^2}+\frac{2e^{-2s}}{s}\]

OpenStudy (zarkon):

\[\mathcal{L}\] \mathcal{L}

OpenStudy (mendicant_bias):

Whoops, typo

OpenStudy (mendicant_bias):

I didn't realize we were able to do that in this editor, thanks!

OpenStudy (anonymous):

You're correct on your answer though, just correct typo of course :)

OpenStudy (anonymous):

Yeah, there's \math__{_} like expressions. I know thats how you do the bold letters for real numbers, complex, etc. \(\mathbb{R}\) = \mathbb{R} etc.

OpenStudy (mendicant_bias):

Oh, okay, cool; thank you!

OpenStudy (anonymous):

Np :3